\begin{thebibliography}{00}
 \bibitem{ALV}
	  Al\`os, E., Le\'on, J.A. and Vives, J., On the short-time
	  behavior of the implied volatility for jump-diffusion models
	  with stochastic volatility. {\it Finance Stoch.} 11, (2007), 
	  571-589.
\bibitem{AlSh}
Al\`os, E. and Shiraya, K., Estimating the Hurst parameter from short
	term volatility swaps. 
{\it Finance Stoch.} 23 (2019),
423-447. 

 \bibitem{BFG}
	  Bayer, C., Friz, P.K. and Gatheral, J., Pricing under rough
	  volatility
	  {\it Quant. Finance} 16:6, (2016),  887-904.
 \bibitem{Friz}
	  Bayer, C. Friz, P.K.,  Gulisashvili, A., Horvath, B. and
	  Stemper B.,
	  Short-time near-the-money skew in  rough fractional volatility
	 models,
{\it Quant. Finance} 19 (2019), 779-798.

\bibitem{Bergomi}
Bergomi, L. Smile dynamics II. {\it Risk}, 10, 67-73, 2005.

\bibitem{CL}
Carr, P. and Lee, R.,
Put-call symmetry : extensions and applications,
{\it Math. Finance} 19 (2009), 523-560.

\bibitem{CW}
Carr, P. and Wu, L., The Finite Moment Log Stable Process and Option
	Pricing,
{\it J. Finance}, LVIII, 2, April 2003.

\bibitem{EFR}
El Euch, O., Fukasawa, M. and Rosenbaum, M., 
The microstructual foundations of leverage
	effect and rough volatility, {\it Finance Stoch.} 22 (2018), 241-280. 

\bibitem{EFGR}
 El Euch, O., Fukasawa, M., Gatheral, J. and Rosenbaum, M.,
Short-term at-the-money
     asymptotics under stochastic volatility models, {\it SIAM
     J. Finan. Math.} 10 (2019), 491-511. 
\bibitem{FO}
Figueroa-L\'opez, J.E. and \'Olafsson, S.,
Short-term asymptotics for the implied volatility skew under a
	stochastic volatility model with L\'evy jumps,
{\it Finance Stoch.} 20 (2016), 973-1020.

 \bibitem{FZ}
	  Forde, M. and Zhang, H., Asymptotics for rough stochastic
	  volatility models,
{\it SIAM J. Finan. Math.}, (2017), 8(1), 114-145. 



  \bibitem{FPS}
	    Fouque, J.P., Papanicolaou, G., Sircar, R. and Solna,
	  K. (2003),
	    Multiscale stochastic volatility asymptotics, {\it
	  Multiscale Model. Simul.} 2, 22-42.

\bibitem{FGP}
Friz, P., Gerhold, S and Pinter, A., 
Option pricing in the moderate deviations regime,
 {\it Math. Finance}, (2018), 28(3), 962-988.

\bibitem{F10} Fukasawa, M., Normalization of implied volatility, (2010),
arXiv:1008.5055

 \bibitem{F11}
	  Fukasawa, M., Asymptotic analysis for stochastic volatility:
	  Martingale expansion. {\it Finance Stoch.} 15, (2011),
	  635-654.
 \bibitem{F2014V}
	  Fukasawa, M. 
 Volatility derivatives and model-free implied leverage, {\it
	 Intern. J. Theoret. Appl. Finance} 17
	 (2014), no.1, 1450002.

 \bibitem{F17}
	  Fukasawa, M., Short-time at-the-money skew and rough
	 fractional
	  volatility, {\it Quant. Finance.} 17:2, (2017), 189-198.


\bibitem{FTW}
Fukasawa, M., Takabatake, T. and Westphal, L., 
Is Volatility Rough ?, (2019),  arXiv:1905.04852


	  \bibitem{GS}
		   Garnier, J. and Solna, K., Correction to
		  Black-Scholes formula
		   due to fractional stochastic volatility, {\it SIAM
		  J. Finan. Math.}, (2017), 8(1) 560-588.
\bibitem{Gas}
Gassiat, P., On the martingale property in the rough Bergomi model,
{\it Electron. Commun. Probab.} 24 (2019), no.33.
		   
\bibitem{Gath}
	 Gatheral, J., {\it The Volatility Surface: A
	 Practioner's Guide}. (2006), John Wiley \& Sons
	 Inc: Hoboken, NJ.
\bibitem{GJR}
	 Gatheral, J., Jaisson, T. and Rosenbaum, M., Volatility is
	 rough.  {\it Quant. Finance.} 18:6, (2018), 933-949.
 \bibitem{GJR2}
	  Guennoun, H., Jacquier, A. and Roome, P., Asymptotic behaviour
	  of the fractional Heston model. 
{\it SIAM J. Finan. Math.}, (2018),  9(3), 1017-1045.

\bibitem{JS}
Jacod, J. and Shiryaev, A., {\it Limit Theorems for Stochastic
	Processes},
2nd ed. (2002), Springer.
\bibitem{JPS}
Jacquier, A., Pakkanen, M.S. and Stone, H.,
Pathwise large deviations for the rough Bergomi model,
{\it J. Appl. Probab.} 55 (2018) 1078-1092.

\bibitem{JR1} Jaisson, T. and Rosenbaum, M.,
Rough fractional diffusions as scaling limits of nearly unstable heavy
	tailed Hawkes processes,
{\it Ann. Appl. Probab.} 26 (2016), 2860-2882.

\bibitem{JR2} Jusselin, P. and Rosenbaum, M.,
No-arbitrage implies power-law market impact and rough volatility,
to appear in {\it Math. Finance} (2019).

\bibitem{Lee}
Lee, R.W., Implied volatility: statics, dynamics, and probabilistic
	 interpretation. {\it Recent Advances in Applied
	 Probability}, 241-268. Springer, Berlin
(2004).

\bibitem{Pigato}
Pigato, P.,
Extreme at-the-money skew in a local volatility model,
{\it Finance Stoch.} 23 (2019), 827-859.

\bibitem{VIX} VIX White Paper, 2019.
 https://www.cboe.com/micro/vix/vixwhite.pdf 

\end{thebibliography}

\documentclass{article}
\usepackage{amsmath,amssymb,pxfonts}
\newtheorem{thm}{Theorem}[section]
\newtheorem{cor}{Corollary}[section]
\newtheorem{rem}{Remark}[section]
\newtheorem{lem}{Lemma}[section]
\title{Volatility has to be rough}
\author{Masaaki Fukasawa\\
{\small Graduate School of Engineering Science, Osaka University}}
\date{}
\begin{document}
 \maketitle
\begin{abstract}
First, we give an asymptotic expansion of 
short-dated at-the-money  implied volatility that refines the preceding
 works and proves in particular that non-rough volatility models are
 inconsistent to  a power law of volatility skew. Second, we show that given a power law of volatility
 skew in an option market, a continuous price dynamics of the underlying asset
 with non-rough volatility admits an arbitrage opportunity.
The volatility therefore has to be rough in a viable market
of the underlying asset of which the volatility skew obeys a power law.
\end{abstract}
\section{Introduction}
It has been almost two decades since 
a power law of volatility skew in option markets 
 was reported \cite{CW,Lee,FPS}.
Denoting $\sigma_{\mathrm{BS}}(k,\theta)$
the Black-Scholes implied volatility with log moneyness $k$ and time to maturity
$\theta > 0$, the power law can be formulated as
\begin{equation*}
 \frac{\sigma_{\mathrm{BS}}(k,\theta) -\sigma_{\mathrm{BS}}(k^\prime,\theta) }
{k-k^\prime} \propto \theta^{H-1/2}
\end{equation*}
for $k \approx 0$ and $k^\prime \approx 0$,  with
 $H\approx 0$, when $\theta \approx 0$.
It is now well known that classical local stochastic volatility models,
where volatility is modeled as a diffusion, are not consistent to the
 power law,
 while some rough volatility models are so \cite{ALV,F11, BFG, F17, GS, FZ,
 GJR2, JPS, EFGR, AlSh, Friz} as well as stable-type discontinuous price models
 \cite{CW,FGP, FO}.
The present article extends the preceding works and 
shows that there is an arbitrage opportunity if volatility is not rough 
given an option market with volatility skew obeying the power law,
under the assumption that the asset price is a positive continuous It\^o
 semimartingale.

In Section 2, we give an asymptotic expansion of short-dated
at-the-money  implied volatility,
which is a refinement of the results in \cite{F17}.
Both the result and proof are much simpler than in \cite{F17} thanks to
choosing the square root of the variance swap fair strike, that is,
VIX~\cite{VIX, Gath}, as the leading term of the expansion, as in
\cite{EFGR}.
Also, we adopt the forward variance framework~\cite{Bergomi,BFG, EFGR} that 
justifies not to consider a time consistency issue treated in \cite{F17}.
In Section~3, we construct the above mentioned arbitrage opportunity.
Some concluding remarks are in Section~4.
All the proofs are given in Appendix.
Throughout the paper, interest rates are assumed to be zero for brevity.

\section{An asymptotic expansion}
\begin{thm}\label{thm1}
Suppose that (the underlying asset price process) $S$ is a positive
 continuous martingale under a measure $Q$ with
$\langle \  \log S \ \rangle$ being absolutely continuous. Let
\begin{equation*}
 V_t = \frac{\mathrm{d}}{\mathrm{d}t} \langle \ \log S \ \rangle_t, \ \ 
v(t) = E[V_t],
\end{equation*}
where $E$ is the expectation under $Q$,
and assume that $v(t)$ is positive and continuous at $t=0$.
If there exists $H \in (0,1/2]$ such that
\begin{equation*}
 \frac{1}{\theta^H}\left(
\frac{V_\theta}{v(\theta)} -1 \right)
\end{equation*}
is uniformly integrable and
 \begin{equation*}
\left(\frac{1}{\sqrt{\theta}}  \left(\frac{S_\theta}{S_0}-1\right),
\frac{1}{\theta^H}\left(
\frac{V_\theta}{v(\theta)} -1 \right)
\right)
 \end{equation*}
converges in law to a two dimensional random variable $(\xi,\eta)$ as
 $\theta \to 0$,
 then
\begin{equation*}
 \sigma_{\mathrm{BS}}\left(z\sqrt{\theta},\theta\right) = 
\sqrt{\bar{v}(\theta)} (1 + \alpha(z)\theta^H)
+  o(\theta^H), \ \text{ as }  \theta \to 0,
\end{equation*}
where $\sigma_{\mathrm{BS}}(k,\theta)$ is the Black-Scholes
 implied volatility as in Introduction at time $t=0$,
\begin{equation*}
\begin{split}
& \bar{v}(\theta) = \frac{1}{\theta} \int_0^\theta v(t)\mathrm{d}t, 
\\
&  \alpha(z) = \frac{1}{2}\int_0^1\int_{-\infty}^\infty
u^HE[\eta | \xi = z \sqrt{u} + \sqrt{v(0)(1-u)}w]\phi(w)\mathrm{d}w \mathrm{d}u,
\end{split}
\end{equation*}
and $\phi$ is the standard normal density.
\end{thm}
\begin{rem}\upshape
 $\sqrt{E[\eta|\xi=x]}$ is a renormalized limit of the
Dupir\'e local volatility.
\end{rem} 
\begin{rem}
\upshape
The function $v$ is called the forward variance curve.
\end{rem}
\begin{rem}
\upshape
The leading term 
 $\sqrt{\bar{v}(\theta)}$ with $\theta = 30$ days corresponds to the VIX.
\end{rem}

\begin{cor}\label{cor1}
Under the condition of Theorem~\ref{thm1},
 if $(\xi,\eta) \sim \mathcal{N}(0,\Sigma)$ with covariance
 matrix $\Sigma = [\Sigma_{ij}]$,
then $\Sigma_{11} = v(0)$, 
$E[\eta | \xi = x] = x \Sigma_{12}/\Sigma_{11}$, and
\begin{equation*}
 \alpha(z) 
= \frac{\Sigma_{12}}{2v(0)} \frac{z}{H+3/2}.
\end{equation*}
In particular, a power law of volatility skew follows: for $\zeta \neq z$,
\begin{equation*}
 \frac{ \sigma_{\mathrm{BS}}\left(z\sqrt{\theta},\theta\right) -
 \sigma_{\mathrm{BS}}\left(\zeta \sqrt{\theta},\theta\right)}
{z\sqrt{\theta} - \zeta \sqrt{\theta} }
\sim \frac{\Sigma_{12}}{\sqrt{v(0)}(2H+3)} \theta^{H-1/2}.
\end{equation*}

\end{cor}

\begin{rem}\upshape
When $V$ is an integrable It\^o semimartingale, 
$v(\theta) = V_0 + O(\theta)$ and
Theorem VIII.3.8 of \cite{JS} verifies
the assumptions of Corollary~\ref{cor1} with $H = 1/2$ and
\begin{equation*}
 (\xi,\eta) \sim \mathcal{N}(0,\Sigma), \ \ 
\Sigma_{11} = \frac{\mathrm{d}}{\mathrm{d}t} 
\langle \log S \rangle_t  \bigg|_{t=0} = v(0), \ \Sigma_{12} = \frac{\mathrm{d}}{\mathrm{d}t} 
\langle \log S, \log V \rangle_t  \bigg|_{t=0}.
\end{equation*}
In particular, for a local volatility model $V_t = \sigma(S_t,t)^2$ with
a smooth function $\sigma$, we have
$v(0) = \sigma(S_0,0)^2$ and $\Sigma_{12} = 2 \sigma(S_0,0) \partial_S\sigma(S_0,0)$.
Thus we conclude the so-called $1/2$ rule :
\begin{equation*}
 \frac{ \sigma_{\mathrm{BS}}\left(z\sqrt{\theta},\theta\right) -
 \sigma_{\mathrm{BS}}\left(\zeta \sqrt{\theta},\theta\right)}
{z\sqrt{\theta} -
 \zeta \sqrt{\theta}  }
\sim \frac{1}{2} \partial_S\sigma(S_0,0).
\end{equation*}
\end{rem}
\begin{rem}\upshape
The martingale property of  the rough Bergomi  model
\begin{equation*}
\begin{split}
 &
 S_t = S_0 \exp\left(\int_0^t\sqrt{V_s}\left[\rho \mathrm{d}W_s
+ \sqrt{1-\rho^2} \mathrm{d}W^\perp_s\right] -\frac{1}{2}\int_0^t V_s\mathrm{d}s\right), 
\\&V_t = v(t) \exp\left(
\int_0^t k(t,s)\mathrm{d}W_s - \frac{1}{2}\int_0^t k(t,s)^2 \mathrm{d}s
\right)
\end{split}
\end{equation*}
with $k(t,s) = \eta |t-s|^{H-1/2}$ was shown by 
\cite{Gas} for $\rho \in [-1,0]$. Using that
\begin{equation*}
\left(H+\frac{1}{2}\right)\theta^{-H-1/2}\int_0^{\theta}k(\theta,s)\mathrm{d}s
 \to \eta, \ \ 
2H \theta^{-2H}\int_0^{\theta}k(\theta,s)^2\mathrm{d}s \to \eta^2 > 0
\end{equation*}
as $\theta \to 0$, 
the assumptions of Corollary~\ref{cor1} are verified with $\Sigma_{12} = \sqrt{v(0)}\rho
\eta/(H + 1/2)$,
and therefore we have a power law of volatility skew
\begin{equation*}
 \frac{ \sigma_{\mathrm{BS}}\left(z\sqrt{\theta},\theta\right) -
 \sigma_{\mathrm{BS}}\left(\zeta
		      \sqrt{\theta},\theta\right)}{z\sqrt{\theta} -
 \zeta \sqrt{\theta}  }
\sim \frac{\rho \ \eta}{(H+1/2)(2H+3)} \theta^{H-1/2}.
\end{equation*}
\end{rem}

\begin{rem}
 \upshape
The model-free implied leverage is defined by \cite{F2014V} 
as the normalized difference 
of the gamma and variance swap fair strikes :
\begin{equation*}
 \lambda(\theta) = \frac{1}{E[\langle \log S \rangle_\theta]}
E\left[
\int_0^\theta \left(\frac{S_t}{S_0}-1\right)\mathrm{d}\langle \log S
 \rangle_t
\right]
 = \frac{\sqrt{\theta}}{ \bar{v}(\theta)}\int_0^1
E[ X^\theta_uV_{\theta u} ]\mathrm{d}u,
\end{equation*}
where $X^\theta$ is defined as (\ref{Xth}) in Appendix.
Under a slightly stronger assumption than in Corollary~\ref{cor1}, namely,
\begin{equation*}
E\left[ \frac{1}{\theta^{H+1/2}}\left(
\frac{S_\theta}{S_0}-1 
\right)\left(\frac{V_\theta}{v(\theta)}-1\right)\right] \to \Sigma_{12},
\end{equation*}
we have
\begin{equation*}
 \theta^{-H}E[X^\theta_u V_{\theta u}]
= E\left[X^\theta_u \theta^{-H}\left(V_{\theta u}
				-v(\theta u)\right)\right]
\to u^{H+1/2} v(0)\Sigma_{12}
\end{equation*}
uniformly in $u \in [0,1]$ and so, a model-free representation of the slope
\begin{equation*}
\frac{\Sigma_{12}}{\sqrt{v(0)}(2H+3)}\theta^{H-1/2}\sim
 \frac{\lambda(\theta)}{2\theta\sqrt{\bar{v}(\theta)}}.
\end{equation*}
\end{rem}
\section{An arbitrage opportunity}
In the previous section we considered the implied volatility at time
$t=0$ and varied the maturity $\theta$.
Here, we fix a maturity $T>0$ instead  and 
consider the implied volatility at time $\tau < T$.
The short-dated asymptotics corresponds to $\tau \uparrow T$.
We start with a lemma that tells about the magnitude of
the Black-Scholes delta hedging error for at-the-money options
when volatility is H\"older continuous.

\begin{lem}
\label{lem1}
Suppose that $S$ is a positive continuous semimartingale with
$\langle \  \log S \ \rangle$ being absolutely continuous. Let
\begin{equation*}
 V_t = \frac{\mathrm{d}}{\mathrm{d}t} \langle \ \log S \ \rangle_t
\end{equation*}
and assume that $V$ is positive and 
$H_0$-H\"older continuous with $H_0 \in (0,1/2]$ a.s.~on $[0,T]$, that is,
\begin{equation*}
\sup_{0 \leq s < t \leq T}\frac{|V_t  - V_s|}{|t-s|^{H_0}} < \infty, \ a.s..
\end{equation*}
Then, for any positive adapted process
$K_\tau$, as $\tau \uparrow T$,
\begin{equation*}
\begin{split}
& 
\left(S_T-\frac{S_\tau^2}{K_\tau}\right)_+
= c_{\mathrm{BS}}(S_\tau,T-\tau) + \int_{\tau}^{T}\frac{\partial
c_{\mathrm{BS}}}{\partial S}(S_t,T-t)\mathrm{d}S_t +
O((T-\tau)^{H_0 + 1/2}), \ a.s.,
\\
& 
(K_\tau-S_T)_+
= p_{\mathrm{BS}}(S_\tau,T-\tau) + \int_{\tau}^{T}\frac{\partial
p_{\mathrm{BS}}}{\partial S}(S_t,T-t)\mathrm{d}S_t +
O((T-\tau)^{H_0 + 1/2}), \ a.s.,
\end{split}
\end{equation*}
and
\begin{equation*}
\begin{split}
&(K_\tau - S_T)_+
- \frac{K_\tau}{S_\tau} \left(
S_T - \frac{S_\tau^2}{K_\tau}
\right)_+ \\ &=  
\int_{\tau}^{T} \left(\frac{\partial
p_{\mathrm{BS}}}{\partial S}(S_t,T-t)
-\frac{K_\tau}{S_\tau}\frac{\partial
c_{\mathrm{BS}}}{\partial S}(S_t,T-t)
\right)\mathrm{d}S_t +
O((T-\tau)^{H_0 + 1/2}), \ a.s.,
\end{split}
\end{equation*}
where $c_{\mathrm{BS}}(S,\theta)$ (resp. $p_{\mathrm{BS}}(S,\theta)$) is
 the Black-Scholes price of 
the call (resp. put) option with the
 underlying asset price $S$, time to maturity $\theta$,  strike price
$S_\tau^2/K_\tau$ (resp.
 $K_\tau$), and
 volatility parameter $\sqrt{V_\tau}$.
\end{lem}

Now we assume a hypothetical option market where call and put options 
with the underlying asset $S$ and 
maturity $T$ are traded at any time $\tau < T$ and for any strike
price $K > 0$. 
Denote by $\sigma_{\mathrm{BS},\tau}(K)$ for
the market implied volatility for the strike price $K$ at time $\tau$.
For $H \in (0,1/2)$,
we say {\it the $H$-power law of negative volatility skew holds} if 
 there exist adapted processes $\sigma_\tau$ and
$\alpha_\tau$ such that
\begin{equation*}
\liminf_{\tau \uparrow T} \sigma_\tau > 0, \ \ 
\limsup_{\tau \uparrow T} \sigma_\tau < \infty, \ \ 
\liminf_{\tau \uparrow T}\alpha_\tau > - \infty, \ \ 
\limsup_{\tau \uparrow T}\alpha_\tau < 0
\end{equation*}
and for any positive adapted process $K_\tau$ with $|K_\tau/S_\tau -1| = O(\sqrt{T-\tau})$,
\begin{equation*}
 \sigma_{\mathrm{BS},\tau}(K_\tau) = \sigma_{\tau} +
  (T-\tau)^{H-1/2}\alpha_\tau \log \frac{K_\tau}{S_\tau}+ o((T-\tau)^H)
\text{ as } \tau \uparrow T.
\end{equation*}
Note that negative volatility skew is typically observed in equity
option markets. What is essential in the following is that 
$\alpha_\tau$ does not change its sign.

Now we construct building blocks of our arbitrage strategy.
Let $\tau_n = T- 1/n$ and 
choose $K_{\tau_n}$ so that $|K_{\tau_n}/S_{\tau_n}-1| = O(n^{-1/2})$ and
\begin{equation*}
\limsup_{n\to \infty} \sqrt{n} \log \frac{K_{\tau_n}}{S_{\tau_n}} < 0.
\end{equation*}
Denote by $\Pi^n$ the P\&L of one unit short of the put option with
strike $K_{\tau_n}$
and $K_{\tau_n}/S_{\tau_n}$ unit long of the call option with strike
$S_{\tau_n}^2/K_{\tau_n}$ with the Black-Scholes delta hedging :
\begin{equation*}
\begin{split}
 \Pi^n =  P_{\tau_n}(K_{\tau_n}) & - \frac{K_{\tau_n}}{S_{\tau_n}}
  C_{\tau_n}\left(\frac{S_{\tau_n}^2}{K_{\tau_n}}\right) +
\int_{\tau_n}^{T} \left(\frac{\partial
p_{\mathrm{BS}}}{\partial S}(S_t,T-t)
-\frac{K_{\tau_n}}{S_{\tau_n}}\frac{\partial
c_{\mathrm{BS}}}{\partial S}(S_t,T-t)
\right)\mathrm{d}S_t  \\
& - (K_{\tau_n} - S_T)_+
+ \frac{K_{\tau_n}}{S_{\tau_n}} \left(
S_T - \frac{S_{\tau_n}^2}{K_{\tau_n}}
\right)_+ ,
\end{split}
\end{equation*}
where $C_\tau(K)$ and $P_\tau(K)$
are respectively the market price of call and put options with strike
$K$ at time $\tau$, and $c_{\mathrm{BS}}$ and $p_{\mathrm{BS}}$
are as in Lemma~\ref{lem1} with $\tau = \tau_n$.

\begin{thm} \label{thm2}
 Suppose the $H$-power law of negative volatility skew holds.
Under the condition of Lemma~\ref{lem1} with $H_0 > H$, 
\begin{equation*}
 \sum_{n=1}^\infty n^{H-1/2}\Pi^n = \infty, \text{ a.s..}
\end{equation*}
\end{thm}
The idea behind Theorem~\ref{thm2} is simple.
If the volatility is $H_0$-H\"older continuous, the Black-Scholes
delta hedging error of the specific option portfolio in the $n$th
building block is only of
$O(n^{-H_0-1/2})$ a.s. by Lemma~\ref{lem1}.
The Black-Scholes price of the portfolio is zero due to the put-call
symmetry~\cite{CL} and the assumed power law of volatility skew 
implies the market price of the portfolio of $O(n^{-H-1/2})$.
That
\begin{equation*}
 \sum_{n=1}^\infty \frac{1}{n} = \infty
\end{equation*}
while
\begin{equation*}
 \sum_{n=1}^\infty \frac{1}{n^{1 + H_0-H}} < \infty
\end{equation*}
enables us to make an almost sure infinite profit.

The implication of Theorem~\ref{thm2} is that in a
viable market,
the volatility cannot have a better H\"older regularity  than $H$, that
is,
it has to be rough.

\section{Concluding remarks}
\begin{rem}\upshape
This paper concludes rough volatility as a consequence of the power law
 in option markets.
The origin of the power law can be explained by a financial practice
 convention.
In FX option markets the convention is to quote prices in terms of the
 implied volatility and tends to quote the same implied volatility for
 the same value of the Black-Scholes delta.
Since the delta is approximately a function of $k/\sqrt{\theta}$, this 
convention makes
$\sigma_{\mathrm{BS}}(z\sqrt{\theta},\theta)$ approximately
 independent of $\theta$, which is nothing but the $H$-power law with $H
 = 0$.
The origin of this convention is not clear.
Naively one may argue that this is
due to the traditional financial engineering that perceives the risk of a
 position only via its delta.
\end{rem}
\begin{rem}\upshape
 The volatility is indeed statistically estimated to be rough;
see~\cite{FTW}.
\end{rem}

\begin{rem}\upshape
 A model-free bound of volatility skew
\begin{equation*}
\left| \frac{\partial \sigma_{\mathrm{BS}}}{\partial k}(0,\theta)\right|
\leq \sqrt{\frac{\pi}{2\theta}}
\end{equation*}
is given in \cite{F10} and shown to be sharp in \cite{Pigato}.
This extreme skew corresponds to the $H$-power law with $H=0$.
Therefore the $H$-power law with $H < 0$ violates no static arbitrage principle in
option markets.
\end{rem}

\begin{rem}
\upshape
 Volatility with regularity $H=0$ can be understood as 
a Gaussian multiplicative chaos. It is however an open question whether there
 exists a continuous-time 
model with both the regularity of $H=0$ and nondegenerate conditional
 skewness that is necessary to recover the power law of volatility skew
 stably in time.
\end{rem}

\begin{rem}\upshape
Derivations of rough volatility as a scaling limit of Hawkes-type market micro
 structure models are given in \cite{JR1, EFR, JR2}.  
In \cite{JR1, EFR}, a heavy-tailed nearly unstable 
self-exciting kernel of order flow 
is the source of the rough volatility.
In \cite{JR2}, such a heavy-tailed kernel is derived via Tauberian
 theorems by assuming
the existing of a scaling limit of market impact functions.
\end{rem}

\begin{rem}\upshape
An inspection of the proof of Lemma~\ref{lem1} reveals that 
the H\"older regularity of volatility only around the maturity $T$ does
 matter.
Therefore a more precise statement of our finding is that 
the volatility has to be rough near the maturities of options.
The volatility has to be rough everywhere 
under a hypothetical framework where vanilla options are traded for any
 strike prices around at-the-money and any maturities.
Note also that our study does not apply any stock price or index whose options
 are not traded.

\end{rem}
\newpage

\begin{appendix}
 \section{ Proof of Theorem~\ref{thm1} }
{\bf Step 1 [An expansion of a rescaled put option price ]}. Denote
\begin{equation}\label{Xth}
 X^\theta_u = \frac{1}{\sqrt{\theta}} 
\left(\frac{S_{\theta u}}{S_0}-1 \right)
\end{equation}
for $u \in [0,1]$. Note that $X^\theta$ is a martingale and with
\begin{equation*}
 \mathrm{d}\langle X^\theta \rangle_u = \left(\frac{S_{\theta
					 u}}{S_0}\right)^2 V_{\theta u}
  \mathrm{d}u = 
(1 + \sqrt{\theta} X^\theta_u)^2 V_{\theta u} \mathrm{d}u.
\end{equation*}
A rescaled put option price can be expressed as
\begin{equation}\label{put}
 \frac{E[(S_0e^{z \sqrt{\theta}}-S_\theta)_+]}{S_0
  \sqrt{\theta}} = E[(\Delta-X^\theta_1)_+], \ \ \Delta = \frac{e^{z
  \sqrt{\theta}} -1}{\sqrt{\theta}}.
\end{equation}
Consider the Bachelier pricing equation with time-dependent variance
\begin{equation*}
 \frac{\partial p}{\partial u} (x,u)
+ \frac{1}{2}v(\theta u) \frac{\partial^2 p}{\partial x^2}(x,u) = 0,\ \ 
p(x,1) = (\Delta - x)_+.
\end{equation*}
The solution  and its derivatives are given
explicitly :
\begin{equation*}
\begin{split}
& p(x,u) = (\Delta-x)\Phi\left(\frac{\Delta-x}{\sqrt{w(1)-w(u)}}\right)
+ \sqrt{w(1)-w(u)}\phi
\left(\frac{\Delta-x}{\sqrt{w(1)-w(u)}}\right), \\
&\frac{\partial p}{\partial x}(x,u) = -
 \Phi\left(\frac{\Delta-x}{\sqrt{w(1)-w(u)}}\right), \\
&\frac{\partial^2 p}{\partial x^2}(x,u) = \frac{1}{\sqrt{w(1)-w(u)}}
 \phi\left(\frac{\Delta-x}{\sqrt{w(1)-w(u)}}\right),
\end{split}
\end{equation*}
where $\Phi$ and $\phi$ are respectively the standard normal
distribution function and the density, and
\begin{equation*}
w(u) = \frac{1}{\theta}\int_0^{\theta u} v(t)\mathrm{d}t.
\end{equation*}
Since the process $X^\theta$ takes values on the interval
$[-\theta^{-1/2}, \infty)$ and the function $p(x,u)$ 
is bounded on  $[-\theta^{-1/2}, \infty) \times [0,1]$ for each $\theta
> 0$,
It\^o's formula, with the aid of a localization argument, gives that
\begin{equation}\label{ito}
\begin{split}
 E[(\Delta-X^\theta_1)_+] &= E[p(X^\theta_1,1)] \\ &= p(0,0) +
\frac{1}{2}E[\int_0^1\frac{\partial^2 p}{\partial x^2}(X^\theta_u,u)
((1 + \sqrt{\theta} X^\theta_u)^2V_{\theta u} - v(\theta u))
\mathrm{d}u].
\end{split}
\end{equation}
By the assumption, 
\begin{equation*}
 \left(X^\theta_u,  \theta^{-H}(
V_{\theta u} - v(\theta u))\right)
\to (X_u,Y_u) :=  (\sqrt{u}\xi, u^H v(0)\eta).
\end{equation*}
We have $\xi \sim \mathcal{N}(0,v(0))$ by the martingale central limit theorem.
Since
\begin{equation*}
 \frac{\partial^2 p}{\partial x^2}(x,u) \to
  \frac{1}{\sqrt{v(0)(1-u)}}\phi\left(
\frac{z-x}{\sqrt{v(0)(1-u)}}
\right)
\end{equation*}
as $\theta \to 0$, we have
\begin{equation*}
 \frac{\partial^2 p}{\partial x^2}(X^\theta_u,u)
\to 
\frac{1}{\sqrt{v(0)(1-u)}}\phi\left(
\frac{z-X_u}{\sqrt{v(0)(1-u)}}\right)
\end{equation*}
in law for each $ u \in [0,1)$.
For any polynomial $q$, there exists a constant $C > 0$ such that
\begin{equation}\label{bound}
\left| q(x)\frac{\partial^2p}{\partial x^2}(x,u) \right| \leq
 \frac{C}{\sqrt{1-u}}.
\end{equation} 
Therefore, the dominated convergence theorem gives that
\begin{equation*}
 \begin{split}
  & 
\theta^{-H}\int_0^1E\left[\frac{\partial^2 p}{\partial
  x^2}(X^\theta_u,u)
(V_{\theta u} - v(\theta u))\right]
\mathrm{d}u \\
& \to 
\int_0^1 E\left[
\frac{1}{\sqrt{v(0)(1-u)}}\phi\left(
\frac{z-X_u}{\sqrt{v(0)(1-u)}}\right)Y_u
\right]\mathrm{d}u
\\ &=  
2\alpha(z)\sqrt{v(0)}\phi\left(\frac{z}{\sqrt{v(0)}}\right)
 \end{split}
\end{equation*}
and that
\begin{equation*}
 \begin{split}
   & 
\int_0^1E\left[\frac{\partial^2 p}{\partial x^2}(X^\theta_u,u)
X^\theta_u V_{\theta u}\right] 
\mathrm{d}u \\
& \to v(0)
\int_0^1 E\left[
\frac{1}{\sqrt{v(0)(1-u)}}\phi\left(
\frac{z-X_u}{\sqrt{v(0)(1-u)}}\right)X_u
\right] \mathrm{d}u
\\ &= 
\frac{z\sqrt{v(0)}}{2}\phi\left(\frac{z}{\sqrt{v(0)}}\right).
 \end{split}
\end{equation*}
From (\ref{put}) and (\ref{ito}), we have then that
\begin{equation} \label{putex}
 \begin{split}
&  \frac{E[(S_0e^{z \sqrt{\theta}}-S_\theta)_+]}{S_0
  \sqrt{\theta}} \\ & = p(0,0) +  
\alpha(z)\sqrt{v(0)}\phi\left(\frac{z}{\sqrt{v(0)}}\right)
  \theta^H +
\frac{z\sqrt{v(0)}}{2}\phi\left(\frac{z}{\sqrt{v(0)}}\right)
  \sqrt{\theta} + o(\theta^H) \\
&=
\Delta \Phi\left(\frac{\Delta}{\sqrt{\bar{v}(\theta)}}\right)
+ \sqrt{\bar{v}(\theta)}
\phi\left(\frac{\Delta}{\sqrt{\bar{v}(\theta)}}\right)
\left(1 + \alpha(z)\theta^H+ \frac{z}{2} \sqrt{\theta}
\right)
  + o(\theta^H).
 \end{split}
\end{equation}
{\bf Step 2 [A comparison with the Black-Scholes model]}.
The Black-Scholes model $\sqrt{V_\theta} \equiv \sigma$, the volatility parameter,
satisfies the assumption with $H = 1/2$ and $\eta = 0$.
Therefore, (\ref{putex}) gives
\begin{equation}\label{BS}
 \frac{P_{\mathrm{BS}}(S_0e^{z\sqrt{\theta}},\theta, \sigma)}{S_0
  \sqrt{\theta}}
=
\Delta \Phi\left(\frac{\Delta}{\sigma}\right)
+ \sigma
\phi\left(\frac{\Delta}{\sigma}\right)
\left(1 + \frac{z}{2} \sqrt{\theta}
\right)
  + o(\theta^{1/2}),
\end{equation}
where $P_{\mathrm{BS}}(K,\theta,\sigma)$ is the Black-Scholes price of
put option with strike $K$, time to maturity $\theta$ and volatility
parameter $\sigma$.
By the Taylor expansion,
\begin{equation*}
 \frac{P_{\mathrm{BS}}(S_0e^{z\sqrt{\theta}},\theta, \sigma +
  a\theta^H)}{S_0 \sqrt{\theta}}
=
\Delta \Phi\left(\frac{\Delta}{\sigma}\right)
+ \sigma
\phi\left(\frac{\Delta}{\sigma}\right)
\left(1 + \frac{z}{2} \sqrt{\theta}
+ \frac{a}{\sigma}\theta^H
\right)
  + o(\theta^H).
\end{equation*}
We can equate this and (\ref{putex}) by setting
\begin{equation*}
\sigma = \sqrt{\bar{v}(\theta)}, \ \ 
 a = \sigma \alpha(z),
\end{equation*}
which implies the result. \hfill////
\\
\section{Proof of Lemma~\ref{lem1}}
Since the Black-Scholes prices $c_{\mathrm{BS}}$ and $p_{\mathrm{BS}}$
satisfy the Black-Scholes equation
\begin{equation*}
\frac{\partial c_{\mathrm{BS}}}{\partial \theta} = \frac{1}{2}V_\tau S^2
\frac{\partial^2 c_{\mathrm{BS}}}{\partial S^2},
\ \ 
\frac{\partial p_{\mathrm{BS}}}{\partial \theta} = \frac{1}{2}V_\tau S^2
\frac{\partial^2 p_{\mathrm{BS}}}{\partial S^2}
\end{equation*}
with
\begin{equation*}
 c_{\mathrm{BS}}(S,0) = \left(S - \frac{S_\tau^2}{K_\tau}\right)_+, \ \ 
 p_{\mathrm{BS}}(S,0) = \left(K_\tau - S\right)_+,
\end{equation*}
It\^o's formula gives
\begin{equation*}
\begin{split}
& 
\left(S_T-\frac{S_\tau^2}{K_\tau}\right)_+
= c_{\mathrm{BS}}(S_\tau,T-\tau) + \int_{\tau}^{T}\frac{\partial
c_{\mathrm{BS}}}{\partial S}(S_t,T-t)\mathrm{d}S_t 
\\ & \hspace*{2cm} +
\frac{1}{2}\int_{\tau}^T (V_t-V_\tau)S_t^2
\frac{\partial^2 c_{\mathrm{BS}}}{\partial S^2}(S_t,T-t)\mathrm{d}t, 
\\
& 
(K_\tau-S_T)_+
= p_{\mathrm{BS}}(S_\tau,T-\tau) + \int_{\tau}^{T}\frac{\partial
p_{\mathrm{BS}}}{\partial S}(S_t,T-t)\mathrm{d}S_t \\
& \hspace*{2cm} +
\frac{1}{2}\int_{\tau}^T (V_t-V_\tau)S_t^2
\frac{\partial^2 p_{\mathrm{BS}}}{\partial S^2}(S_t,T-t)\mathrm{d}t.
\end{split}
\end{equation*}
Since $|V_t-V_\tau| \leq C|t-\tau|^{H_0}$ for some finite random
variable $C$ by the assumption and 
\begin{equation*}
 \left|\frac{\partial^2 c_{\mathrm{BS}}}{\partial S^2}(S_t,T-t)
\right| \vee
 \left|\frac{\partial^2 p_{\mathrm{BS}}}{\partial S^2}(S_t,T-t)
\right| \leq \frac{1}{\sqrt{2\pi V_\tau(T-t)} \inf_{t \in [\tau,T]}S_t},
\end{equation*}
we obtain the first two equations. 
The last equation follows from the first two with aid of 
the put-call symmetry~\cite{CL}:
\begin{equation*}
 p_{\mathrm{BS}}(S_\tau,T-\tau) = \frac{K_\tau}{S_\tau}c_{\mathrm{BS}}(S_\tau,T-\tau).
\end{equation*}
\hfill{////}

\section{Proof of Theorem~\ref{thm2}}
Let
\begin{equation*}
Z_n =\sqrt{n} \log \frac{K_{\tau_n}}{S_{\tau_n}}.
\end{equation*}
Then, $\liminf_{n\to \infty} Z_n > -\infty$,
$\limsup_{n \to \infty} Z_n < 0$ and
\begin{equation*}
\begin{split}
& \sigma_{\mathrm{BS},\tau_n}(K_{\tau_n}) = \sigma_{\tau_n} +
  n^{-H}\alpha_{\tau_n} Z_n + o(n^{-H}),\\
& \sigma_{\mathrm{BS},\tau_n}(S_{\tau_n}^2/K_{\tau_n}) = \sigma_{\tau_n} -  n^{-H}\alpha_{\tau_n} Z_n + o(n^{-H})
\end{split}
\end{equation*}
by the assumed power law.
The Taylor expansion of the Black-Scholes price with respect to
the volatility parameter gives
\begin{equation*}
 P_{\tau_n}(K_{\tau_n}) = p_{\mathrm{BS}}+
  \frac{\partial p_{\mathrm{BS}}}{\partial \sigma}
  n^{-H}\alpha_{\tau_n}Z_n + o(n^{-H})
\end{equation*}
and
\begin{equation*}
 C_{\tau_n}(S_{\tau_n}^2/K_{\tau_n}) = c_{\mathrm{BS}}-
  \frac{\partial c_{\mathrm{BS}}}{\partial \sigma}
  n^{-H}\alpha_{\tau_n}Z_n + o(n^{-H}),
\end{equation*}
where $p_{\mathrm{BS}}$ and $c_{\mathrm{BS}}$ are the Black-Scholes
prices with volatility parameter $\sigma_{\tau_n}$ of,
respectively,
put option with strike $K_{\tau_n}$ and
call option with strike $S_{\tau_n}^2/K_{\tau_n}$.
By the put-call symmetry~\cite{CL} of the Black-Scholes prices,
\begin{equation*}
 P_{\tau_n}(K_{\tau_n}) -
\frac{K_{\tau_n}}{S_{\tau_n}} C_{\tau_n}(S_{\tau_n}^2/K_{\tau_n}) = 
 \left( \frac{\partial p_{\mathrm{BS}}}{\partial \sigma} + \frac{K_{\tau_n}}{S_{\tau_n}}\frac{\partial c_{\mathrm{BS}}}{\partial \sigma}  \right)
  n^{-H}\alpha_{\tau_n}Z_n + o(n^{-H}).
\end{equation*}
Note that  $\liminf_{n \to \infty} Z_n > -\infty$ ensures 
\begin{equation*}
\liminf_{n\to \infty} 
\sqrt{n} \left( \frac{\partial p_{\mathrm{BS}}}{\partial \sigma} +
	  \frac{K_{\tau_n}}{S_{\tau_n}}\frac{\partial c_{\mathrm{BS}}}{\partial \sigma}  \right) > 0.
 \end{equation*}
Further, we have $\liminf_{n\to \infty} \alpha_{\tau_n}Z_n > 0$ and so,
\begin{equation*}
 \sum_{n=1}^\infty
n^{H-1/2}\left( P_{\tau_n}(K_{\tau_n}) -
\frac{K_{\tau_n}}{S_{\tau_n}} C_{\tau_n}(S_{\tau_n}^2/K_{\tau_n})\right)  = \infty.
\end{equation*}
The result then follows from Lemma~\ref{lem1}. \hfill////

\end{appendix}

\end{document}